# Contrastive Representations for Temporal Reasoning

### v1

@@ Appendix D Evaluation Details
We evaluate all networks on $1000$ problem instances per environment. For Rubik’s Cube, each instance is a cube scrambled using 1000 moves. For 15-Puzzle, Lights Out, and Digit Jumper, evaluation boards are sampled randomly. For Sokoban, we follow the same instance generation procedure as described by Czechowski et al., (2021).
@@ Appendix F Generalization to Temporal Reasoning in Non-Combinatorial Domains
To investigate whether CRTR also identifies temporal features in non-combinatorial domains, we apply it to a dataset of robotic manipulation trajectories (the Adroit dataset from D4RL Fu et al., (2020)). Those tasks require using a high-dimensional robotic hand to perform fine motor activities, and are designed to test fine motor control and long-horizon planning. We quantify representation quality by measuring the predicted distance from each state in a trajectory to the final state in a trajectory. Specifically, we look at the rank correlation between the time step and predicted distance, with a correlation of 1 indicating that the learned representations are highly predictive of the temporal distance from each state to the final state.
@@ Appendix H Mutual Information Analysis
To estimate the conditional mutual information, we use NPEET package, which implements the method proposed in (Kraskov et al., , 2004) that uses k-nearest neighbours for entropy estimation. We conduct the analysis using trajectories collected from the Sokoban or Digit Jumper environment, utilizing all transitions within these trajectories (>45k>45k transitions for Sokoban and >20k>20k for Digit Jumper). The variables used in the experiment are defined as follows:

### v2

@@ Appendix D Evaluation Details
We evaluate all networks on $1000$ problem instances per environment. For Rubik’s Cube, each instance is a cube scrambled using 1000 moves. For 15-Puzzle, Lights Out, and Digit Jumper, evaluation boards are sampled randomly. For Sokoban, we follow the same instance generation procedure as described by Czechowski et al., [14].
@@ Appendix F Generalization to Temporal Reasoning in Non-Combinatorial Domains
To investigate whether CRTR also identifies temporal features in non-combinatorial domains, we apply it to a dataset of robotic manipulation trajectories (the Adroit dataset from D4RL [23]). Those tasks require using a high-dimensional robotic hand to perform fine motor activities, and are designed to test fine motor control and long-horizon planning. We quantify representation quality by measuring the predicted distance from each state in a trajectory to the final state in a trajectory. Specifically, we look at the rank correlation between the time step and predicted distance, with a correlation of 1 indicating that the learned representations are highly predictive of the temporal distance from each state to the final state.
@@ Appendix H Mutual Information Analysis
To estimate the conditional mutual information, we use NPEET package, which implements the method proposed in [37] that uses k-nearest neighbours for entropy estimation. We conduct the analysis using trajectories collected from the Sokoban or Digit Jumper environment, utilizing all transitions within these trajectories (>45k>45k transitions for Sokoban and >20k>20k for Digit Jumper). The variables used in the experiment are defined as follows: